OpenStudy (anonymous):

Paraceratherium was the biggest land mammal there has ever been. It lived about 35 million years ago and was 8 m tall and 11 m long. It looked like a giant rhinoceros but had a long neck like a giraffe. About how tall was Paraceratherium in feet? Use the conversion factor 1 ft. = 0.3048 m. Round to the nearest tenth where necessary. a. 26.7 ft. c. 42.6 ft. b. 36.7 ft. d. 19.9 ft. gIVING A MEDAL

OpenStudy (e.mccormick):

Know how to set up a ratio?

OpenStudy (e.mccormick):

Personal Messages are not a way to harass other users. Like all content on OpenStudy, personal messages are subject to the Code of Conduct you agreed to for using this site. http://openstudy.com/code-of-conduct Read the section under "Don't be offensive, inappropriate, or creepy!" Each thing you are dealing with can be put as \(\dfrac{\text{Feet}}{\text{Meters}}\). Once you have two ratios (fractions) set them equal to each other: \(\dfrac{\text{Feet Standard}}{\text{Meters Standard}}=\dfrac{\text{Feet Unknown}}{\text{Meters Given}}\) Then solve for the one you don't have a number for.